OpenStudy (anonymous):

Find the consecutive integers m and n such that the given number is between m and n, or state the the given number is not a real number. 1. -√115

OpenStudy (campbell_st):

think about \[-\sqrt{121}\] \[-\sqrt{115}\] \[-\sqrt{100}\] this may help

OpenStudy (anonymous):

Still confusing

OpenStudy (anonymous):

So the answer will be the square root of this

OpenStudy (campbell_st):

\[-\sqrt{121} = -11\] \[-\sqrt{115}\] \[-\sqrt{10} = -10\] then the number is between - 11 and - 10.... so it is a real number

OpenStudy (anonymous):

Thanks a lot